OpenStudy (anonymous):

Natasha rides her bike (at a constant speed) for 3 hours, helped by a wind of 3 miles per hour. Pedaling at the same rate, the trip back against the wind takes 9 hours. Find find the total round trip distance she traveled

OpenStudy (anonymous):

t=s/v 3= s/v1 +v2 and 9 = s/v1-v2 you can do it silly ;)

OpenStudy (anonymous):

what does any of that mean????? o.O

OpenStudy (anonymous):

lol v1 is velocity of bike, v2 is velocity of wind

OpenStudy (anonymous):

can't you just give me the answer?

OpenStudy (anonymous):

nope that is against the law

OpenStudy (anonymous):

Well, you're no help at all.

OpenStudy (anonymous):

thank you ;)